The weekly revenue R (in dollars) earned by a computer repair service is given by R= -12p^2+372p, where p is the price charged per service hour (in dollars).
a) find the weekly revenues for prices per service hour of $12, $16, and $20.
b) find the price that will yield a maximum weekly revenue.
c) what is the maximum weekly revenue?

Answer by ewatrrr(24785)

R= -12p^2+372p
a) R($12) = $2736, R($16) = $2880, R($20) = $2640
|
R= -12p^2+372p = -12(p- 31/2) + 12(31/2)^2 = -12(p-31/2) +2883
b) $15.50
c) $2883